OpenStudy (anonymous):

I really really need this question answered: Consider the following function. y= 1/(x-8)^3 Which two functions below make up a composite function form for y? y=1/u u=1/x u=x-8 y=u^3 y=u^-3 u=x^3

OpenStudy (anonymous):

if we assume u=(x-8)^3 then y=1/u

OpenStudy (anonymous):

yes but i can only choose 2 of the functions..

OpenStudy (anonymous):

if u=(x-8), then y=u^-3

OpenStudy (anonymous):

y=u^-3 , y=1/u